The gravitational attraction between two rocks of equal mass in deep space is 16.0 N. If both rocks double in mass and the distance separating the two rocks doubles , what is the resulting gravitational force between them ?

(A) 8.00 N

(B) 12.0 N

(C ) 16.0 N

(D ) 32.0 N


1
Expert's answer
2021-04-23T11:08:37-0400

Solution.

"F=16.0N;"

"F=G\dfrac{m_1m_2}{r^2};"

"F_1=G\dfrac{2m_1\sdot 2m_2}{(2r)^2}=G\dfrac{4m_1m_2}{4r^2}=G\dfrac{m_1m_2}{r^2};"

"F_1=F=16.0N;"

Answer: "(C) 16.0 N."